# Brain Organoid Computing – an Overview

### v1

@@ Abstract
The aim of this paper is to give an overview of brain organoid computing, its characteristics, challenges, as well as possible advantages for future applications in the field of artificial intelligence. An important part is the extensive bibliography covering all relevant aspects and questions on this topic.
@@ I Introduction
Recent advances in biomedical engineering practices have led to increased research on the use of in vitro brain cultures known as brain organoids to test responses to stimuli and investigate neural structures [mulder_beginners_2023]. Despite the mainly medical focus [koo_past_2019] of this research, a novel field of using these neural structures for computational tasks was created [smirnova_organoid_2023, jordan_open_2024].
Research on using biological tissue for computational purposes has so far largely been focused on using biological components to create circuits that mimic current silicon-based computers. However, this method of biocomputing forgoes the use of the full potential of cells in this context [grozinger_pathways_2019]. Yet, novel research in the field of leveraging brain organoids for computing has enabled early applications, such as FinalSpark’s Neuroplatform, that relies on human brain organoids and facilitates remote access for universities and institutions to experiment with their organic processors [jordan_open_2024]. The use of brain organoids and more primitive neuronal cultures for computing related tasks is a field with a comparatively short history. Since the advent of induced pluripotent stem cells (iPSCs), research on such use cases has ranged from neuronal cultures performing basic image processing [ruaro_toward_2005], moving avatars in virtual environments [bakkum_spatio-temporal_2008], to blind source separation [isomura_cultured_2015].
Following the creation of the first brain organoids, subsequent research on the use of in vitro neurons for computing has generated quite some interest [smirnova_organoid_2023]. Despite the still speculative capabilities of brain organoids, researchers have begun to define and investigate how these can be further exploited for computing. Accordingly, current research focuses on developing platforms and architectures to enable novel utilizations of brain organoids in computing contexts, as well as improving the very organoids forming the substrate of this technology [smirnova_organoid_2023, jordan_open_2024, johns_hopkins_university_first_2024].
The organoids used in experiments are generated using iPSCs, placed in a suitable medium, and connected to a multielectrode array (MEA) [govindan_mass_2021]. The MEAs allow to read out electrical signals from the organoids as well as to electrically stimulate them in order to influence their behavior [bakkum_spatio-temporal_2008, kagan_vitro_2022]. Using these methods researchers have been able to modify the organoids’ activity by providing feedback by means of electrical signals [bakkum_spatio-temporal_2008, jordan_open_2024]. Thus, such brain organoids are addressable and respond to stimuli, opening the possibility of influencing their activity and steering their processing capabilities for computational tasks.
@@ II Opportunities and Challenges
With traditional silicon processors approaching physical limits with respect to integration density, clock frequencies, etc., processors are iteratively providing less performance gain per generation [leiserson_theres_2020] than in previous decades. This means that the main process driving hardware performance improvements is becoming harder to maintain and may soon stop delivering the gains that have been crucial to the advancement of technology today. Furthermore, increasing integration densities, despite the gradual deceleration in their progress, leads to quicker degradation of logic gates, lowering overall performance over time, which also limits further advances in this direction [cao_future_2023].
Additionally, modern silicon-based processing units are impractical for research in fields like AI due to their high power requirements. Current top-of-the line machine learning models already require electric power at a Megawatt scale [smirnova_organoid_2023]. Estimates show that data centers used for big data and AI applications accounted for 460 terawatt-hours in 2022 alone [iea_electricity_2024], which is even higher than France’s energy consumption in 2022 [eia_electricity_2024]. In addition, today’s predominantly von Neumann processors are limited with respect to the achievable amount of parallelism [schuman_opportunities_2022]. They also struggle to deal with decision making based on complex and uncertain datasets [smirnova_organoid_2023], both being advantageous for AI tasks [basu_organoid_2024]. Therefore, current processors use vast amounts of electricity for tasks they are not optimized for, while traditional methods of improving processing power are reaching fundamental limits. Consequently, researchers have been exploring alternative methods of computing such as using organic tissue that has been shown to exhibit intelligent behavior processing information, stimuli and contexts, mimicking real biological systems.
@@ II-A1 Adaptivity
Brain organoids can adapt to stimuli and given tasks, successively increasing their competence for the current task and have been shown to do so within a range of activities. This ability to modify and strengthen behaviors and neural structures in response to experiences and stimuli is known as plasticity [jeon_distinctive_2023]. In vitro cultures of neurons have, for example, learned to play Pong [kagan_vitro_2022], learned to conduct speech recognition and conduct simple computations [cai_brain_2023]. In said examples the systems continuously improved their performance. Experiments like these showcase the capability of in vitro neural structures to plastically learn from experiences and change their behavior to optimize performance.
This is in contrast with traditional computing efforts, where the underlying hardware architecture is rigid and only software and its parametrization can be changed dynamically. The adaptability of these cultures enables an architectural plasticity that can complement the adaptability of the software, enabling flexible and efficient computing pathways [basu_organoid_2024].
This advantage not only applies to novel tasks and problems, but also aids traditional machine learning tasks by minimizing learning epochs [cai_brain_2023, jeon_distinctive_2023]. Biological neural networks (BNNs) and their inherent plasticity could therefore enable systems to learn like biological brains do and progressively improve their performance in given tasks, improving efficiency. Furthermore, their plasticity allows them to predictively process inputs. This stems from the deeply interconnected nature of BNNs and their evolutionarily advantageous attempts to minimize surprises to the system by anticipating and extrapolating future activity based on past inputs [luczak_neurons_2022, jeon_distinctive_2023].
Meanwhile, contemporary artificial neural networks (ANNs) predominantly rely on interpolation within the bounds of their training data, making them less adaptable to unpredictable environments and data outside of their training datasets [bonnasse-gahot_interpolation_2022, taheri_generalization_2024]. Experiments comparing BNNs and ANNs highlight this, with BNNs significantly outperforming ANNs in pattern recognition in randomized shapes [fleuret_comparing_2011].
This basic capability is further improced by their inherent plasticity. Accordingly, BNNs can adapt and respond quickly to new stimuli and while the network improves over time to effectively deal with a wider range of stimuli [basu_organoid_2024, stoop_note_2021, luczak_neurons_2022]. These capabilites are well demonstrated by BNNs adapting to perform blind source separation [isomura_cultured_2015], as well as predicting auditory stimuli [luczak_neurons_2022]. In both cases the networks were exposed to training sets for short periods of time.
@@ II-A2 Parallelism
One of the main disadvantages of traditional von Neumann or Harvard central processing unit (CPU) architectures is the inherent sequential processing of instructions. This is in stark contrast to BNNs which exhibit full parallelism due to the absence of any central control logic, memory, etc. [schuman_opportunities_2022]. This characteristic has the potential to solve multiple tasks in parallel [ruaro_toward_2005, basu_organoid_2024], thus decreasing processing times and therefore creating a vastly more efficient system than one that is based solely on sequential (digital) processing.
A further advantage of brain organoid computing lies in its unconventional signaling and information communication method. Neurons emit electrical signals when firing, so-called spikes or action potentials [sharf_functional_2022] in contrast to bits used in conventional computers. [^1] It should be noted that such spiking behavior occurs in most but not all biological neurons. Spikes transmit more information than just an on / off state as they encode information in their amplitude, waveform, as well as their duration, and the overall spiking patterns [sharf_functional_2022]. Spikes an not only triggered by stimuli but can also occur spontaneously [sharf_functional_2022].
These multidimensional signals allow for a comparatively high amount of information encoded in transmitted signals [jeon_distinctive_2023]. Additionally, spikes are processed by neurons on an event-driven basis, meaning that neurons perform work mainly when spikes are present in the network [schuman_opportunities_2022].
@@ II-A3 Efficiency
Accordingly, this results in a highly energy efficient processing [liu_unleashing_2024]. Using brain organoids for computational tasks could eventually result in massive efficiency gains for complex information processing systems. Given the current and projected energy requirements of modern data centers, this is a key benefit.
Hewlett Packard Enterprise’s Frontier supercomputer requires around 24.6 megawatts of electricity to achieve an average performance of 1.353 EFLOPS ($10^{18}$ floating point operations per second) [TOP500_2024], which, although not easily, can be compared to the raw computational power of a human brain. Several sources estimate a human brain capable of performing between $10^{15}$[kurzweil_singularity_2005] and $10^{18}$[madhavan_brain-inspired_2023] FLOPS using roughly 20 Watts [smirnova_organoid_2023]. This makes a brain more power efficient by a factor ranging from $10^{3}$ to $10^{6}$ – a remarkable disparity. Biological brains are fundamentally more energy efficient than any modern digital system.
Training a large language model like OpenAI’s GPT-3 is estimated to require around 1,300 MWh of electricity [luccioni_estimating_2022], which is equivalent to the electricity consumption of more than 800 households in the Europaen Union during one year [eurostat_electricity_2024]. The exceptional energy efficiency of BNNs results mainly from their inherently parallel operation and their spiking behavior with associated event driven computation. Using biological tissue allows for highly efficient computing architectures. This might lead to further breakthroughs in AI and associated fields [basu_organoid_2024], despite using less electricity than current systems.
In addition to these advantages brain organoids for use in computing could be very economical. Current costs for generating a brain organoid containing roughly $10^{5}$ (the number of connections is highly variable and can not specified here) cells can be as low as €0.36, with the maintenance cost for one year being around €0.66 per organoid [govindan_mass_2021].
Nevertheless, there is no consensus on the potential costs of mature brain organoid-based computing, as there currently is no broad real-world implementation. In fact, due to the advanced and multidisciplinary technology and know-how required for such a fully-fledged system, some researchers argue that this technology may be significantly more costly when increasing its complexity [basu_organoid_2024].
@@ II-A4 Mortal Computing
Brain organoids can also be used to create self-organizing systems whose hardware and software are inseparable and unique to the system, known as mortal computers [ororbia_mortal_2023] (this might also be true for analog electronic implementation of such networks due to unavoidable parameter variations inherent to a particular implementation). These proposed systems rely on the self-organization, adaption, and homeostasis of the chosen substrate and low-level processing [ororbia_mortal_2023, hinton_forward-forward_2022]. The concept of mortal computing is based on the notion of embodiment, where a system’s computational power arises from the complex interplay of its physical structure and processing capabilities, such as in the case of brain organoids.
Such systems have been proposed to be highly efficient [hinton_forward-forward_2022] and adaptive, [ororbia_mortal_2023], characteristics that have also been used to describe brain organoids when used for computing [smirnova_organoid_2023]. Mortal computers have been proposed to be a pathway to the creation of AGI and towards advances in neuromorphic systems and computational models [ororbia_mortal_2023]. With their plasticity and parallel processing, brain organoids offer a potential substrate for such systems, enabling more adaptive, efficient, and evolving computational structures.
@@ II-B1 Lifespan
The biggest challenges of using brain organoids for computational purposes lie in their biological nature. First of all, brain organoids have a limited lifespan ranging from around 100 days [jordan_open_2024] to 15 months and more, depending on the method of maintenance and what their application [govindan_mass_2021].
Due to the lack of circulatory systems and the dependence on a neuronal medium to sustain them, the tissue eventually starts necrosis and stops developing [smirnova_promise_2024]. Current brain organoid lifespans represent only 25 percent of the average life expectancy of modern supercomputers at maximum [rojas_analyzing_2019]. Consequently, computing systems using brain organoids as their processing base would have a significantly shorter operational lifespan compared to traditional computers. Solving this problem is one of the most important tasks in this field [sun_generation_2022].
Additionally, the organoids are sensitive to environmental changes and experience cell stress and necrosis even under optimal conditions, thus complicating research and application considerably [zabolocki_brainphys_2020, kim_application_2023].
@@ II-B2 Reproducibility
Brain organoids grown from stem cells exhibit significant variability in cellular composition and structure due to their very nature and various (uncontrollable) developmental factors. Even under controlled conditions, organoids differentiate spontaneously, reducing model comparability and leading to errors, uncertain outcomes – all significant challenges [qian_brain_2019]. Their sensitivity to factors such as cell-culture conditions, generation methods, sustaining ingredients, etc., further compounds these issues [poli_experimental_2019, qian_brain_2019]. This unpredictability undermines practical applications, as unreliable systems are unsuitable for critical tasks requiring accuracy [mittal_survey_2016].
In addition to these uncertainties the spikes emitted and transmitted by the neurons induce noise, further reducing the reproducibility of results. Moreover, neurons can fire spontaneously as part of their normal behavior [sharf_functional_2022]. However, these spontaneous spikes typically are not a problem [bakkum_spatio-temporal_2008, smirnova_organoid_2023] and can even be used as a baseline for activity in experiments [cai_brain_2023].
Finally, the issue of reproducibility can slow down future progress and research in this field. All in all, brain organoids’ inherent spontaneity proves to be a difficult challenge to overcome in research and future of uses in a computational context.
@@ II-B3 Processing Power
Although neurons are comparatively slow and unreliable processing units, they benefit from large numbers of them being interconnected densly [ruaro_toward_2005, luczak_neurons_2022]. However, current brain organoids are still quite limited in the number of neurons available and their respective interconnectivity.
Though fully developed human brains are comparable in raw processing power to modern supercomputers, current brain organoids are only comparable to early embryonic brains with up to $\approx 10^{5}$ cells, whereas a fully developed human brain contains around $10^{11}$ cells with about $10^{14}$ synapses (connections) [smirnova_organoid_2023].
Generally, higher intelligence and cognition correlates with the complexity of the brain structures [goriounova_large_2018]. Accordingly, human brain organoids are currently severely limited in their capacity to replicate advanced processing as found in human brains. This will continue at least in the near future due to the absence of a circulatory system, resulting in oxygen deprivation, stress, limited maturation, and thus reduced lifespan [kim_application_2023], massively impeding their processing potential.

### v2

@@ Abstract
The aim of this paper is to give an overview of brain organoid computing, its characteristics, challenges, as well as possible advantages for future applications in the field of artificial intelligence. An important part is the extensive bibliography covering all relevant aspects and questions on this topic. Brain organoids—three-dimensional in vitro neural structures derived from human stem cells—have recently garnered attention not only in medical research but also as potential substrates for unconventional computing. Their biological nature allows them to exhibit learning behavior, plasticity, and parallel information processing, making them fundamentally different from traditional silicon-based systems. This opens up new perspectives on how intelligent systems might be designed in the future. Using brain organoids for computing presents a possible pathway towards more adaptive, energy-efficient, and biologically inspired forms of AI. However, challenges persist, particularly regarding lifespan, interfacing, reproducibility, and ethical concerns regarding the use of human-derived tissue. This paper aims to provide a foundational understanding for researchers exploring the convergence of human biology and computation.
@@ I Introduction
Recent advances in biomedical engineering practices have led to increased research on the use of in vitro brain cultures known as brain organoids to test responses to stimuli and investigate neural structures [1]. Despite the mainly medical focus [2] of this research, a novel field of using these neural structures for computational tasks was created [3, 4].
Research on using biological tissue for computational purposes has so far largely been focused on using biological components to create circuits that mimic current silicon-based computers. However, this method of biocomputing forgoes the use of the full potential of cells in this context [5]. Yet, novel research in the field of leveraging brain organoids for computing has enabled early applications, such as FinalSpark’s Neuroplatform, that relies on human brain organoids and facilitates remote access for universities and institutions to experiment with their organic processors [4]. The use of brain organoids and more primitive neuronal cultures for computing-related tasks is a field with a comparatively short history. Since the advent of induced pluripotent stem cells (iPSCs), research on such use cases has ranged from neuronal cultures performing basic image processing [6], moving avatars in virtual environments [7], to blind source separation [8].
Following the creation of the first brain organoids, subsequent research on the use of in vitro neurons for computing has generated quite some interest [3]. Despite the still speculative capabilities of brain organoids, researchers have begun to define and investigate how these can be further exploited for computing. Accordingly, current research focuses on developing platforms and architectures to enable novel utilizations of brain organoids in computing contexts, as well as improving the very organoids forming the substrate of this technology [3, 4, 9].
The organoids used in experiments are generated using iPSCs, placed in a suitable medium, and connected to a multielectrode array (MEA) [10]. The MEAs allow researchers to read out electrical signals from the organoids as well as to electrically stimulate them in order to influence their behavior [7, 11]. Using these methods researchers have been able to modify the organoids’ activity by providing feedback by means of electrical signals [7, 4]. Thus, such brain organoids are addressable and respond to stimuli, opening the possibility of influencing their activity and steering their processing capabilities for computational tasks.
@@ II Opportunities and Challenges
With traditional silicon processors approaching physical limits with respect to integration density, clock frequencies, etc., processors are iteratively providing less performance gain per generation [12] than in previous decades. This means that the main process driving hardware performance improvements is becoming harder to maintain and may soon stop delivering the gains that have been crucial to the advancement of technology today. Furthermore, increasing integration densities, despite the gradual deceleration in their progress, leads to quicker degradation of logic gates, lowering overall performance over time, which also limits further advances in this direction [13].
Additionally, modern silicon-based processing units are impractical for research in fields like AI due to their high power requirements. Current top-of-the line machine learning models already require electric power at a Megawatt scale [3]. Estimates show that data centers used for big data and AI applications accounted for 460 terawatt-hours in 2022 alone [14], which is even higher than France’s energy consumption in 2022 [15]. In addition, today’s predominantly von Neumann processors are limited with respect to the achievable amount of parallelism [16]. They also struggle to deal with decision-making based on complex and uncertain datasets [3], both being advantageous for AI tasks [17]. Therefore, current processors use vast amounts of electricity for tasks they are not optimized for, while traditional methods of improving processing power are reaching fundamental limits. Consequently, researchers have been exploring alternative methods of computing such as using organic tissue that has been shown to exhibit intelligent behavior processing information, stimuli and contexts, mimicking real biological systems.
@@ II-A1 Adaptivity
Brain organoids can adapt to stimuli and given tasks, successively increasing their competence for the current task and have been shown to do so within a range of activities. This ability to modify and strengthen behaviors and neural structures in response to experiences and stimuli is known as plasticity [18]. In vitro cultures of neurons have, for example, learned to play Pong [11], learned to conduct speech recognition and simple computations [19]. In said examples the systems continuously improved their performance. Experiments like these showcase the capability of in vitro neural structures to plastically learn from experiences and change their behavior to optimize performance.
This is in contrast with traditional computing efforts, where the underlying hardware architecture is rigid and only software and its parametrization can be changed dynamically. The adaptability of these cultures enables an architectural plasticity that can complement the adaptability of the software, enabling flexible and efficient computing pathways [17].
This advantage not only applies to novel tasks and problems, but could also help with traditional machine learning tasks by minimizing learning epochs [19, 18]. Biological neural networks (BNNs) and their inherent plasticity could therefore enable systems to learn as biological brains do and progressively improve their performance in given tasks, improving efficiency. Furthermore, their plasticity allows them to process inputs predictively. This stems from the deeply interconnected nature of BNNs and their evolutionarily advantageous attempts to minimize surprises to the system by anticipating and extrapolating future activity based on past inputs [20, 18].
Meanwhile, contemporary artificial neural networks (ANNs) predominantly rely on interpolation within the bounds of their training data, making them less adaptable to unpredictable environments and data outside of their training datasets [21, 22]. Experiments comparing BNNs and ANNs highlight this, with BNNs significantly outperforming ANNs in pattern recognition in randomized shapes [23].
This basic capability is further improved by their inherent plasticity. Accordingly, BNNs can adapt and respond quickly to new stimuli while the network improves over time to effectively deal with a wider range of stimuli [17, 24, 20]. These capabilities are well demonstrated by BNNs adapting to perform blind source separation [8], as well as predicting auditory stimuli [20]. In both cases the networks were exposed to training sets for short periods of time.
@@ II-A2 Parallelism
One of the main disadvantages of traditional von Neumann or Harvard central processing unit (CPU) architectures is the inherent sequential processing of instructions. This is in stark contrast to BNNs which exhibit full parallelism due to the absence of any central control logic, memory, etc. [16]. This characteristic has the potential to solve multiple tasks in parallel [6, 17], thus decreasing processing times and therefore creating a vastly more efficient system than one that is based solely on sequential (digital) processing.
A further advantage of brain organoid computing lies in its unconventional signaling and information communication method. Neurons emit electrical signals when firing, so-called spikes or action potentials [25] in contrast to bits used in conventional computers. [^1] It should be noted that such spiking behavior occurs in most but not all biological neurons. Spikes transmit more information than just an on / off state as they encode information in their amplitude, waveform, as well as their duration, and the overall spiking patterns [25]. Spikes are not only triggered by stimuli but can also occur spontaneously [25].
These multidimensional signals allow for a comparatively high amount of information encoded in transmitted signals [18]. Additionally, spikes are processed by neurons on an event-driven basis, meaning that neurons perform work mainly when spikes are present in the network [16].
@@ II-A3 Efficiency
Accordingly, this results in a highly energy efficient processing [26]. Using brain organoids for computational tasks could eventually result in massive efficiency gains for complex information processing systems. Given the current and projected energy requirements of modern data centers, this is a key benefit.
Hewlett Packard Enterprise’s Frontier supercomputer requires around 24.6 megawatts of electricity to achieve an average performance of 1.353 EFLOPS ($10^{18}$ floating point operations per second) [27], which, although not easily, can be compared to the raw computational power of a human brain. Several sources estimate a human brain capable of performing between $10^{15}$[28] and $10^{18}$[29] FLOPS using roughly 20 Watts [3]. This makes a brain more power efficient by a factor ranging from $10^{3}$ to $10^{6}$ – a remarkable disparity. Biological brains are fundamentally more energy efficient than any modern digital system.
Training a large language model like OpenAI’s GPT-3 is estimated to require around 1,300 MWh of electricity [30], which is equivalent to the electricity consumption of more than 800 households in the European Union during one year [31]. The exceptional energy efficiency of BNNs results mainly from their inherently parallel operation and their spiking behavior with associated event driven computation. Using biological tissue allows for highly efficient computing architectures. This might lead to further breakthroughs in AI and associated fields [17], despite using less electricity than current systems.
In addition to these advantages brain organoids for use in computing could be very economical. Current costs for generating a brain organoid containing roughly $10^{5}$ (the number of connections is highly variable and can not specified here) cells can be as low as €0.36, with the maintenance cost for one year being around €0.66 per organoid [10].
Nevertheless, there is no consensus on the potential costs of mature brain organoid-based computing, as there currently is no broad real-world implementation. In fact, due to the advanced and multidisciplinary technology and know-how required for such a fully-fledged system, some researchers argue that this technology may be significantly more costly when increasing its complexity [17].
@@ II-A4 Mortal Computing
Brain organoids can also be used to create self-organizing systems whose hardware and software are inseparable and unique to the system, known as mortal computers [32] (this might also be true for analog electronic implementation of such networks due to unavoidable parameter variations inherent to a particular implementation). These proposed systems rely on the self-organization, adaption, and homeostasis of the chosen substrate and low-level processing [32, 33]. The concept of mortal computing is based on the notion of embodiment, where a system’s computational power arises from the complex interplay of its physical structure and processing capabilities, such as in the case of brain organoids.
Such systems have been proposed to be highly efficient [33] and adaptive, [32], characteristics that have also been used to describe brain organoids when used for computing [3]. Mortal computers have been proposed to be a pathway to the creation of AGI and towards advances in neuromorphic systems and computational models [32]. With their plasticity and parallel processing, brain organoids offer a potential substrate for such systems, enabling more adaptive, efficient, and evolving computational structures.
@@ II-B1 Lifespan
The biggest challenges of using brain organoids for computational purposes lie in their biological nature. First of all, brain organoids have a limited lifespan ranging from around 100 days [4] to 15 months and more, depending on the method of maintenance and intended application [10].
Due to the lack of circulatory systems and the dependence on a neuronal medium to sustain them, the tissue eventually undergoes necrosis and stops developing [34]. Current brain organoid lifespans represent only 25 percent of the average life expectancy of modern supercomputers at maximum [35]. Consequently, computing systems using brain organoids as their processing base would have a significantly shorter operational lifespan compared to traditional computers. Solving this problem is one of the most important tasks in this field [36].
Additionally, the organoids are sensitive to environmental changes and experience cell stress and necrosis even under optimal conditions, thus complicating research and application considerably [37, 38].
@@ II-B2 Reproducibility
Brain organoids grown from stem cells exhibit significant variability in cellular composition and structure due to their very nature and various (uncontrollable) developmental factors. Even under controlled conditions, organoids differentiate spontaneously, reducing model comparability and leading to errors, uncertain outcomes – all significant challenges [39]. Their sensitivity to factors such as cell-culture conditions, generation methods, sustaining ingredients, etc., further compounds these issues [40, 39]. This unpredictability undermines practical applications, as unreliable systems are unsuitable for critical tasks requiring accuracy [41].
In addition to these uncertainties the spikes emitted and transmitted by the neurons induce noise, further reducing the reproducibility of results. Moreover, neurons can fire spontaneously as part of their normal behavior [25]. However, these spontaneous spikes typically are not a problem [7, 3] and can even be used as a baseline for activity in experiments [19].
Finally, the issue of reproducibility can slow down future progress and research in this field. Brain organoids’ inherent spontaneity proves to be a difficult challenge to overcome in research and future of uses in a computational context.
@@ II-B3 Processing Power
Although neurons are comparatively slow and unreliable processing units, they benefit from large numbers of them being interconnected densely [6, 20]. However, current brain organoids are still quite limited in the number of neurons available and their respective interconnectivity.
Though fully developed human brains are comparable in raw processing power to modern supercomputers, current brain organoids are only comparable to early embryonic brains with up to $\approx 10^{5}$ cells, whereas a fully developed human brain contains around $10^{11}$ cells with about $10^{14}$ synapses (connections) [3].
Generally, higher intelligence and cognition correlates with the complexity of the brain structures [42]. Accordingly, human brain organoids are currently severely limited in their capacity to replicate advanced processing as found in human brains. This will continue at least in the near future due to the absence of a circulatory system, resulting in oxygen deprivation, stress, limited maturation, and thus reduced lifespan [38], massively impeding their processing potential.